How many molecules are in 5 moles of ammonia, NH3?

1
Expert's answer
2022-03-12T01:58:55-0500

n(NH3) = 5 moles;

NA(Avogadro constant) = 6.02 * 1023 mol-1;

n = N/NA;

N(NH3) = n(NH3) * NA = 5 * 6.02 * 1023 = 30.1* 1023 = 3.01 * 1024 molecules.

Answer: 3.01 * 1024.